# How to Bridge Structural and Temporal Heterogeneity in Link Prediction? A Contrastive Method

### v1

@@ 3.2. Temporal Information Modeling Layer
Time-level Heterogeneity Differentiation Modeling. To bridge the nonuniformity of temporal sequences between long- and short-term sequence learning spaces, we propose a time-level contrastive learning method to approach the latent long-term and short-term latent sequential representations, respectively.
@@
(16) $\mathcal{L}^{S+}_{Time}=\sum_{a=1}^{\left|V^{T}\right|}\log\frac{\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{S}\cdot\mathbf{u}_{a}^{L}\right)/\tau\right)}{\sum_{b\neq a}\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{S}\cdot\mathbf{u}_{b}^{S}\right)/\tau\right)},$
@@ 3.3. Output Layer
Then the total loss $\mathcal{L}_{total}$ is formulated by weighting three specific types of losses: the main cross-entropy loss $\mathcal{L}_{{main}}$, node-level heterogeneity differentiation loss denoted by $\mathcal{L}_{N}=\mathcal{L}_{{Node}}^{{+}}-\mathcal{L}_{\text{Node }}^{{-}}$, edge-level heterogeneity differentiation loss expressed as $\mathcal{L}_{E}=\mathcal{L}_{{Edge}}^{{+}}-\mathcal{L}_{{Edge}}^{{-}}$, and time-level heterogeneity differentiation loss represented by $\mathcal{L}_{T}=\mathcal{L}^{L+}_{{Time}}+\mathcal{L}^{S+}_{{Time}}$.
@@
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ represent the learnable weighting factors employed to balance three losses. The procedure of our CLP is outlined in Algorithm 1. We define $d$, $N$, and $T$ as the node embedding dimension, the number of node and snapshot, respectively. The time complexity of our CLP is $O(tNd^{2})$.
Algorithm 1 Link Prediction with CLP.
0:  Final node representation embedding: $\mathbf{u}_{a}$.
1:  for Each train iteration do
2:     Parameters initialization.
3:     for $\mathcal{G}^{t}$ in $\mathcal{G}$ do
4:        Construct subgraph $\mathcal{G}^{rt}$ based on the edge type $r\in R$.
5:        for Every subgraph $\mathcal{G}^{rt}$ associated with edge type $r$ do
9:           Calculate the node-level loss $\mathcal{L}_{Node}^{+}$ and $\mathcal{L}_{Node}^{-}$ via Eq. (5) and Eq. (6);
10:        end for
11:        Calculate the edge-level weight $\delta_{a}^{rt}$ for any node $a$ via Eq. (7) and Eq. (8);
14:        Calculate the edge-level loss $\mathcal{L}_{Edge}^{+}$ and $\mathcal{L}_{Edge}^{-}$ via Eq. (11) and Eq. (12);
15:     end for
18:     Calculate the time-level loss $\mathcal{L}_{Time}^{L+}$ and $\mathcal{L}_{Time}^{S+}$ via Eq. (15) and Eq. (16) ;
19:     Calculate the total loss $\mathcal{L}_{total}$ via Eq. (17) and Eq. (18).
20:  end for
21:  return  $\mathbf{u}_{a}$
@@ 4. Experiments and Analysis
  • Q1:How does CLP’s performance compare to state-of-the-art models?
@@
  • Q3: How does adjusting hyperparameters affect the CLP’s performance?
@@
4.1. Experimental Setup
4.1.1. Datasets
We perform experiments on four application scenarios, i.e., Math-overflow, Taobao, OGBN-MAG, and COVID-19, to verify the universality and effectiveness of our proposed CLP. Details of four datasets are presented in Table 2.
Math-overflow  [^1] http://snap.stanford.edu/data/sx-mathoverflow.html.: Math-overflow serves as an interactive website for mathematics enthusiasts and professionals. It adopts a forum-style model for communication and interaction, fostering an online community of expert mathematicians. This dataset comprises 2350 days of user interactions, which can be categorized into three types: question and answer exchanges, question and comment discussions, and answer and comment engagements. For analytical purposes in our experiments, we divide this dataset into 11 snapshots using a time window of 124 days.
Taobao (Du et al., 2019): Taobao is characterized by three types of node: user (U), product (I), and topic (T), along with three types of links: U-I, U-T, and I-T, which signify users’ interactions on cloud-themed products in the Taobao App from April 1 to May 31, 2008. For our experimentation, this dataset is divided into five snapshots using a time window of 12 days.
OGBN-MAG (Hu et al., 2020b): OGBN-MAG is a subset of Microsoft Academic Graph (MAG), which contains four node types (papers, authors, institutions, and research areas) and four types of relationships (authors affiliated with institutions, authors writing papers, papers citing other papers, and papers associated with research areas as their topics). In this experiment, a THN is extracted from OGBN-MAG for the period from January 1st to 10th, 2010. Time is divided into time slots, each containing 1000 edges of each type to form the OGBN-MAG dataset used in our experiment.
COVID-19  [^2] https://coronavirus.1point3acres.com/en: COVID-19, sourced from 1point3acres, provides state-level and county-level daily case reports, including confirmed cases, new cases, deaths, and recovered cases. We select the daily new COVID-19 cases as our THN data. This dataset consists of two node types: state and county) and three relationships: one administrative affiliation (state includes county) and two geospatial relationships (state adjacent to state, county adjacent to county). In our experimental study, the constructed THN covers the period from May 1 to 21, 2020 and comprises 21 time snapshots. Each snapshot restricts each relationship type to a maximum of 2000 edges.
After the time window is partitioned, further data cleansing is necessitated. In our model and the baselines, representation vectors of nodes are derived from the initial $t$ network snapshots. However, it is not feasible to obtain representation vectors for nodes that fail to appear in the $(t+1)$-th network snapshots, rendering link prediction for these nodes impossible. Consequently, nodes that newly appear, along with their corresponding links in the $(t+1)$-th snapshot, are eliminated. Additionally, the negative links for the training set are collected from links that are absent in the initial $t$ snapshots. For the test set, negative links are sampled from those not present in the $(t+1)$-th snapshot.
Table 2. Statistics of Datasets.
Dataset Math-overflow Taobao OGBN-MAG COVID-19
#Nodes 24818 29475 17269 2165
#Edges 506550 63367 40000 131649
Node Type 1 3 4 2
Edge Type 3 3 4 3
#Snapshots 11 5 10 21
Time Granularity Second Second Year Day
4.1.2. Baselines
We evaluate our CLP against eight baselines categorized into four groups: Static Homogeneous, Static Heterogeneous, Dynamic Homogeneous, and Dynamic Heterogeneous approaches. These baselines encompass both traditional and advanced link prediction models that are highly relevant to our research. In addition, we provide detailed descriptions of the fundamental elements of the baseline models and our CLP in Table 3.
(1) Static Homogeneous approaches:
  • SEAL (Zhang and Chen, 2018) extracts the local graph for the target link and learns the local graph features to estimate the probability of that link’s existence.
  • VGNAE (Ahn and Kim, 2021) integrates variational inference, GCNs, and normalization techniques to effectively learn probabilistic node embeddings from graph-structured data.
(2) Static Heterogeneous approaches:
  • Metapath2Vec (Dong et al., 2017) leverages metapath-guided random walks combined with the skip-gram model to learn low-dimensional embeddings for nodes in heterogeneous information networks, capturing both structural and semantic relationships within the network.
  • GATNE (Cen et al., 2019) learns node representations in attributed multiplex networks. It successfully addresses the challenges of integrating multiple types of relationships and node attributes into a unified representation.
(3) Dynamic Homogeneous approaches:
  • TGAT (Xu et al., 2020) presents an inductive representation learning approach for temporal graphs, combining temporal encoding and GNNs to capture the dynamic nature of such graphs, which can be generalized to new nodes and future graph snapshots.
  • TDGNN (Qu et al., 2020) designs a continuous-time link prediction method for dynamic graphs, leveraging temporal encodings and attention mechanisms to enhance the predictive capabilities of GNNs.
(4) Dynamic Heterogeneous approaches:
  • THAN (Li et al., 2023) leverages memory mechanisms and transformer architectures to capture intricate temporal and structural information of temporal heterogeneous graphs.
  • THGAT (Zhang et al., 2023a) combines neighborhood type modeling, neighborhood information aggregation, and time encoding technique to achieve accurate node representations.
(5) Our models:
Table 3. Comparisons of Baseline Models and Our Proposed Approaches.
Model Key Components
Static Homogeneous SEAL (Zhang and Chen, 2018) Node Labeling, Subgraph Extraction, GNN Graph Learning
Baselines VGNAE (Ahn and Kim, 2021) Graph Autoencoder, Variational Inference, GCN Graph Learning
Static Heterogeneous Metapath2Vec (Dong et al., 2017) Metapath-guided Random Walks, Skip-gram Modeling
GATNE (Cen et al., 2019) Attributed Multiplex Networks, GCN Graph Learning
Dynamic Homogeneous TGAT (Xu et al., 2020) Graph Attention Network, Self-attention Mechanism
TDGNN (Qu et al., 2020) GCN Graph Learning, Temporal Aggregator, Edge Aggregator
Dynamic Heterogeneous THAN (Li et al., 2023) Dynamic Transfer Matrix, Self-attention Mechanism
THGAT (Zhang et al., 2023a) Neighborhood Type Modeling and Aggregation, Temporal Dynamics Integration
Our model CLP Node-level Feature Modeling, Edge-level Feature Modeling, Temporal Information Modeling
CLP-N Edge-level Feature Modeling, Temporal Information Modeling
CLP-E Node-level Feature Modeling, Temporal Information Modeling
CLP-T Node-level Feature Modeling, Edge-level Feature Modeling
4.1.3. Implementation Details
We have made the source code and datasets publicly available at https://github.com/tayer915/CLP.git. During the training process, our CLP is executed with a batch size of 1024. We employ an early-stopping strategy halting training when the Average Precision (AP) metric ceases to increase for $5$ consecutive epochs. The learning rate $lr$ is set at 1e-4. We employ Adam as the optimizer of our CLP model. The hyper-parameters are carefully optimized following a grid search. Specifically, the dimension $d$ of node embeddings is explored within the range of $\left\{8,16,32,64,128\right\}$. Moreover, we search for the optimal values of the balance coefficients $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ in the range of $1e\left\{-4,-6,-8,-9,-10\right\}$; the number of attention heads $h$ in the range of $\left\{1,2,4,8,16\right\}$; and the temparature coefficient $\tau$ in the closed interval of $[0.04,0.12]$ with a step size of $0.02$. By default, after fine-tuning, we adopt the following hyperparameter settings, wherein the optimal values of $d,\lambda_{\cdot},h,\text{ and }\tau$ are set to $0.132,1e-8,4,\text{ and }$ 0.1, respectively.
The node representations are derived from the first $T$ network snapshots $\{\mathcal{G}^{1},\mathcal{G}^{2},\cdots,\mathcal{G}^{T}\}$. The links in the subsequent network snapshot $\mathcal{G}^{T+1}$ serve as the evaluation set. Moreover, 20% of the links in this evaluation set are randomly assigned as the validation set, another 20% as the positive training set, and the remaining 60% as the positive test set. Simultaneously, several non-existent negative links are randomly sampled to form negative training and test sets.
4.1.4. Evaluation Metrics
We employ widely-adopted Area Under the Curve (AUC) and AP as evaluation metrics (Li et al., 2023; Xu et al., 2020). AUC represents the area under the Receiver Operating Characteristic (ROC) curve, which plots the false positive rate on the x-axis and the true positive rate on the y-axis. AP refers to the area under the Precision-Recall curve, with recall on the horizontal axis and precision on the vertical axis. Values of AUC and AP closer to 1 indicate superior model performance.
4.2. Performance and Analysis
@@
SEAL (Zhang and Chen, 2018) 63.33 63.68 51.26 55.11 69.66 74.74 60.12 63.62
VGNAE (Ahn and Kim, 2021) 62.65 57.06 58.40 54.80 56.85 53.72 72.89 65.59
Metapath2Vec (Dong et al., 2017) 64.62 72.30 50.20 66.48 62.32 74.11 54.66 67.32
GATNE (Cen et al., 2019) 71.53 73.87 71.32 82.71 68.01 77.63 71.15 66.30
TGAT (Xu et al., 2020) 66.46 60.25 77.49 71.38 61.72 63.06 61.44 64.62
TDGNN (Qu et al., 2020) 83.09 74.69 71.33 73.84 78.94 70.66 67.28 65.77
THAN (Li et al., 2023) 89.23 89.05 65.22 65.51 74.63 71.88 74.07 70.49
THGAT (Zhang et al., 2023a) 77.27 80.88 80.04 84.30 85.97 86.42 64.88 67.62
@@ 4.2.3. Effect of Parameters (for Q3)
In response to Q3, we conduct a sensitivity analysis of the key parameters of our model on Taobao and OGBN-MAG. These parameters include the dimension of vector representation ($d$), the number of multi-heads ($h$), loss weights ($\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$) and the temperature coefficient ($\tau$).
@@
(3) Loss weights ($\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$): Adjusting the weights for node-, edge-, and time-level contrastive heterogeneity differentiation loss is critical for balanced optimization during model training. Precise tuning of these parameters ensures that each component contributes proportionally to its role in processing heterogeneous data, enhancing overall model performance. The optimal prediction performance is achieved when $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are all set to $81e-$ 8, as depicted in Figure 6. This setting notably minimizes errors in the final output by effectively balancing the emphasis on different types of losses, customizing our CLP to specific tasks and environments, pivotal in achieving high performance in dynamic heterogeneous data-driven scenarios.
@@
(4) Temperature coefficient ($\tau$): $\tau$ modulates the sensitivity to sample similarities during training, impacting overall model efficacy. Specifically, optimal $\tau$ value enables the contrastive loss function to accurately differentiate between similar and dissimilar samples. As shown in Figure 7, the value of $0.1$ yields the most effective outcomes, significantly optimizing contrastive learning performance within our CLP. This setting encourages the model’s focus on meaningful variances among samples, thereby improving both accuracy and robustness in real-world scenarios. Adjusting sensitivity through $\tau$ is instrumental in refining the model’s ability to generalize from training data to unseen data, a critical aspect for heterogeneous and dynamic environments.
5. Related Work
In this section, we present a comprehensive overview of researches related to our work, encompassing three distinct branches: Heterogeneous Network Embedding approaches, Temporal Network Embedding approaches, and Contrastive Learning approaches.
5.1. Heterogeneous Network Embedding Approaches
Heterogeneous network embedding approaches aim to capture the structural and semantic data in complex networks characterized by diverse types of nodes and edges, presenting unique challenges for link prediction (Wang et al., 2022). Studies often address structural and semantic heterogeneity by constructing meta-paths (Dong et al., 2017; Fu et al., 2017; Qiu et al., 2018; Wang et al., 2019; Zhao et al., 2023b). Specifically, Metapath2vec (Dong et al., 2017) and HIN2Vec (Fu et al., 2017) employ random walks guided by meta-paths to sample heterogeneous information, enabling effective link prediction in heterogeneous networks. Moreover, Metapath2vec++ (Qiu et al., 2018) considers proximity measures and meta-path weighting schemes, while LPMPA (Zhao et al., 2023b) considers the node pairs as the learning target by mapping the network into multiple semantic graphs using various meta-paths. Building on the concept of meta-paths, mg2vec (Zhang et al., 2020) incorporates meta-graphs into the process of node embedding learning. However, these methods may ignore node and edge attributes, leading to a focus on methods that explore such attributes extensively. Specifically, HGT (Hu et al., 2020a) adopts hierarchical attention mechanisms to aggregate information from diverse neighbor types, thereby enhancing link prediction accuracy across varied domains. TALP (Li et al., 2020) captures an $n$-tuple representation through a two-layer graph attention architecture and considers the role of local type information in aligning user nodes for anchor link prediction. Other approaches like THGNN (Xu et al., 2021) and TH-SLP(Peng et al., 2023) emphasize topic-aware heterogeneous neural network that establish multi-dimensional representations for link prediction. Likewise, LHGNN (Nguyen et al., 2023) develops representations of latent heterogeneous graphs through node and path-level embedding, and PaGE-Link (Zhang et al., 2023d) focuses on elucidating GNN explanations in heterogeneous networks, thereby enhancing link prediction capabilities. These approaches demonstrate the effectiveness of message passing in capturing complex relational patterns in heterogeneous networks. However, they fail to assign independent learnable parameters to edges, thereby constraining the model’s capacity to acquire specific edge features. The aforementioned heterogeneous network embedding methods have substantially advanced link prediction. However, their effectiveness largely depends on the accurate integration of diverse data types.
5.2. Temporal Network Embedding Approaches
Temporal graph learning approaches hold vast potential for addressing the dynamic nature of real-world networks and can be categorized into two categories: sequential and graph models. Sequential models (Hao et al., 2020; Chen et al., 2019; Jiang et al., 2023; Zhang et al., 2023c) divide the THNs into a series of temporal snapshots and employ RNNs and attention mechanisms to learn snapshot dependencies. For instance, Hao et al. (Hao et al., 2020) utilize a GRU to simulate the progression of node vectors and predict the future node representations. E-LSTM-D (Chen et al., 2019) combines LSTM with an encoder-decoder architecture for dynamic network link prediction. DNformer (Jiang et al., 2023) proposes a transfer learning framework for temporal link prediction that equates the self-attention module used in link sequences to traditional graph embedding techniques. AMCNet (Zhang et al., 2023c) enhances dynamic link prediction through multi-scale representation, attention mechanisms, and a co-evolving model. on the other hand, graph models (Lei et al., 2019; Yang et al., 2019; Pareja et al., 2020) are better suited for handling variable-length sequence data and adapting to the fluctuating nature of time-series data compared to sequence models. Specifically, GCN-GAN (Lei et al., 2019) and NetworkGAN (Yang et al., 2019) concentrate on weighted dynamic networks using Graph Convolutional Networks (GCNs) to investigate the topological properties of each snapshot. EvolveGCN (Pareja et al., 2020) dynamically updates the GCN parameters as the graph evolves over time, effectively capturing the changing nature of relationships within the network. TGAT(Xu et al., 2020) presents a temporal graph attention approach for dynamic network embedding, which integrates self-attention mechanisms to learn the features of temporal links. ComGCN (Pham et al., 2021) captures network characteristics at microscopic, mesoscopic, and macroscopic levels and employs k-layered GCN to capture the macroscopic characteristics of the network. MetaDyGNN (Yang et al., 2022) targets the few-shot link prediction scenarios using a dynamic GNN aligned with meta-learning principles. HTGNN (Fan et al., 2022) designs multiple layers of heterogeneous temporal aggregation through intra-relation, inter-relation, and across-time aggregation modules. GSNOP (Luo et al., 2023) leverages GNNs and neural ordinary differential equations for link prediction in dynamic, sparse graphs. TGGDN (Huang and Lei, 2023) integrates temporal dynamics, graph diffusion, and group-aware modeling to enhance methodological efficiency. DHGAS (Zhang et al., 2023e) explores optimal GNN architectures through a dynamic neural architecture search framework. Temporal graph learning approaches have shown significant advantages in handling diverse aspects of temporal networks, particularly in capturing and predicting temporal dynamics. However, there are ongoing challenges in their implementation, such as the demands on computational resources and the capabilities for real-time processing.

### v3

@@ 3.2. Temporal Information Modeling Layer
Time-level Heterogeneity Differentiation Modeling. To bridge the nonuniformity of temporal sequences between long- and short-term sequence learning spaces, we propose a time-level contrastive learning method to approach the latent long-term and short-term latent sequential representations, respectively:
@@
$\mathcal{L}^{S+}_{Time}=\sum_{a=1}^{\left|V^{T}\right|}\log\frac{\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{S}\cdot\mathbf{u}_{a}^{L}\right)/\tau\right)}{\sum_{b\neq a}\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{S}\cdot\mathbf{u}_{b}^{S}\right)/\tau\right)},$
@@ 3.3. Output Layer
[…]
@@
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ represent the learnable weighting factors employed to balance three losses. The procedure of our CLP is outlined in Algorithm 1 in the Appendix. We define $d$, NN, and $T$ as the node embedding dimension, the number of nodes, and snapshots, respectively. The time complexity of our CLP is $O(TNd^{2})$.
@@ 4. Experiments and Analysis
  • Q1: How does CLP perform compared to state-of-the-art models?
@@
  • Q3: How do hyperparameters affect the CLP’s performance?
@@
SEAL (zhang2018link) 63.33 63.68 51.26 55.11 69.66 74.74 60.12 63.62
VGNAE (ahn2021variational) 62.65 57.06 58.40 54.80 56.85 53.72 72.89 65.59
Metapath2Vec (dong2017metapath2vec) 64.62 72.30 50.20 66.48 62.32 74.11 54.66 67.32
GATNE (cen2019representation) 71.53 73.87 71.32 82.71 68.01 77.63 71.15 66.30
TGAT (XuRKKA20) 66.46 60.25 77.49 71.38 61.72 63.06 61.44 64.62
TDGNN (qu2020continuous) 83.09 74.69 71.33 73.84 78.94 70.66 67.28 65.77
THAN (li2023memory) 89.23 89.05 65.22 65.51 74.63 71.88 74.07 70.49
THGAT (zhang2023dynamic) 77.27 80.88 80.04 84.30 85.97 86.42 64.88 67.62
@@
4.1. Experimental Setup
4.1.1. Datasets and Metrics
We perform experiments on four application scenarios, i.e., Math-overflow, Taobao, OGBN-MAG, and COVID-19, to verify the universality and effectiveness of our proposed CLP. We employ widely-adopted Area Under the Curve (AUC) and AP as evaluation metrics (li2023memory; XuRKKA20). The specifics and processing details of four datasets are presented in $\S$ B.2 of the Appendix.
4.1.2. Baselines
We evaluate our CLP against eight baselines categorized into four groups: (1) Static Homogeneous(SEAL (zhang2018link), VGNAE (ahn2021variational)), (2) Static Heterogeneous (Metapath2Vec (dong2017metapath2vec), GATNE (cen2019representation)), (3) Dynamic Homogeneous (TGAT (XuRKKA20), TDGNN (qu2020continuous)), and (4) Dynamic Heterogeneous (THAN (li2023memory), THGAT (zhang2023dynamic)). Further detailed elaboration on baseline models and our model is available in $\S$ B.3 of the Appendix.
4.1.3. Implementation Details
We have made the source code and datasets publicly available at https://anonymous.4open.science/r/CLP-7E66. During the training process, our CLP is executed with a batch size of 1024. We employ an early-stopping strategy halting training when the Average Precision (AP) metric ceases to increase for $5$ consecutive epochs. The learning rate $lr$ is set at 1e-4. We employ Adam as the optimizer of our CLP model. The hyper-parameters are carefully optimized following a grid search. More details can be found in $\S$ B.1 of the Appendix.
4.2. Performance and Analysis
@@ 4.2.3. Effect of Parameters (for Q3)
In response to Q3, we conduct a sensitivity analysis of the key parameters of our model on Taobao and OGBN-MAG. These parameters include the dimension of vector representation ($d$), the number of multi-heads ($h$), loss weights ($\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$) and the temperature coefficient ($\tau$).
@@
(3) Loss weights ($\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$): Adjusting the weights for node-, edge-, and time-level contrastive heterogeneity differentiation loss is critical for balanced optimization during model training. Precise tuning of these parameters ensures that each component contributes proportionally to its role in processing heterogeneous data, enhancing overall model performance. The optimal prediction performance is achieved when $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are all set to 1e31e-3, as depicted in Figure 6. This setting notably minimizes errors in the final output by effectively balancing the emphasis on different types of losses, customizing our CLP to link prediction task, pivotal in achieving high performance in dynamic heterogeneous scenarios.
(4) Temperature coefficient ($\tau$): $\tau$ modulates the sensitivity to sample similarities during training, impacting overall model efficacy. Specifically, optimal $\tau$ value enables the contrastive loss function to accurately differentiate between similar and dissimilar samples. As shown in Figure 7, the value of $0.1$ yields the most effective outcomes, significantly optimizing contrastive learning performance within our CLP. This setting encourages the model’s focus on meaningful variances among samples, thereby improving both accuracy and robustness in real-world scenarios.
5. Related Work
5.1. Heterogeneous Network Embedding Approaches
Heterogeneous network embedding approaches focus on capturing structural and semantic features in complex networks with diverse node and edge types to improve link prediction accuracy (wang2022survey). Some studies (dong2017metapath2vec; fu2017hin2vec; wang2019heterogeneous; qiu2018network; zhao2023link; zhang2020mg2vec) employ meta-paths and meta-graphs for efficient sampling of heterogeneous information. However, these methods may overlook node and edge attributes. To overcome such limits, some researches, including (hu2020heterogeneous; li2020type; nguyen2023link; zhang2023page), pay their attention to aggregating diverse neighbor information, while others (xu2021topic; peng2023th) focus on the topic-aware node attributes. However, these methods rely on the accurate integration of diverse data types, and lack the in-depth inspection of data distribution.
5.2. Temporal Network Embedding Approaches
Temporal graph learning approaches can be divided into sequential and graph models to handle the dynamic nature of real-world networks. Sequential models (hao2020dynamic; chen2019lstm; jiang2023dnformer; zhang2023attentional) divide the THNs into a series of temporal snapshots and employ RNNs (hao2020dynamic; chen2019lstm) and attention mechanisms (jiang2023dnformer; zhang2023attentional) to learn snapshot dependencies. Meanwhile, graph models (lei2019gcn; yang2019advanced; pareja2020evolvegcn) are better suited for handling variable-length sequence data and adapting to the fluctuating nature of time-series data compared to sequence models. Specifically, models including (lei2019gcn; yang2019advanced; pareja2020evolvegcn; XuRKKA20) leverage GNNs to capture evolving relationships and topological properties within dynamic networks. Other researches (pham2021comgcn; yang2022few; fan2022heterogeneous; luo2023graph; huang2023temporal; zhang2023dynamic1) further enhance dynamic network embedding through various aggregation and modeling strategies. However, despite the advantages of these approaches in capturing temporal dynamics, challenges remain in terms of computational demands.